Two hundred tickets were sold for a baseball game, which amounted to $840. Student tickets cost $3, and adult tickets cost $5.
A. Let x be the nuber of student tickets sold and y be the number of adult tickets sold. Writ a system of linear equations whose solution give the nuber of each type of ticket sold.
B. Use the method of substitution to solve the system.

Answer by jorel1380(3719)

x+y=200
3x+5y=840
3x=(840-5y)
3x+3y=600
840-5y+3y=600
240=2y
y=120
x=80
There were 80 student tickets sold, and 120 adult tickets sold.
80(3)+120(5)=240+600=840..